Y^2+x^2p^2-2xyp=4/p^2
1
Expert's answer
2020-12-23T16:50:34-0500
"p^2x^2-2pxy+y^2=\dfrac{4}{p^2}""(y-px)^2=\dfrac{4}{p^2}"

Or

"y=px\pm\dfrac{2}{p}"

Each equation is in Clairaut's form. Hence changing "p" to the arbitrary constant "c," the required solution is


"(y-cx)^2=\dfrac{4}{c^2}, c \text{ is an arbitrary constant}"